Calculate the values ​​of U1, U3, U4, U5, I1, I2, I3, I4, RT and IT in the circuit below.

Ry=32 Iz? Rz=682 U=6V 1=? Uz=? Rs=? 1=? Ry=622 V=? Ri=352 HE U=? U= 24 V
Answer #1

By using ohm's law and node voltage method, we can find the above given terms.

0 7 Given circuit is, R₂ = 32 I R3=662 U2=60 Iz Uz = ? Rag = 7 U4=? R4=652 NI; = ? U5 = ? YI4=? Ri=32 + U,=? U-24 V we can m2 I, = I4 ; because they 24-4.5 3+3+1 = 1.5A 13 13 I4 =15A 7 Iz 24.5 = 6 and - = 4.5 R4 6 ,مملح * UI * U4 Ry 1 R34=302 Serie(3 a. Ry=162 To find If Jy ó- T Now, replacing Rt with all resistors now the given circuit will + Ry = 162 U=24v IT 24 - :. I